OpenStudy (epicteatime):

Discuss why the degree of the resulting denominator did not change from your expression’s degree.

OpenStudy (epicteatime):

what does this mean...

OpenStudy (epicteatime):

\[\frac{ 2x }{ (x-5) }\times \frac{ (x+3) }{ 6 }\]

OpenStudy (epicteatime):

@Hero

OpenStudy (epicteatime):

\[\frac{ 2x(x+3) }{ 6(x-5) }\]

OpenStudy (epicteatime):

@satellite73

OpenStudy (epicteatime):

yeah i had another problem...

OpenStudy (anonymous):

a rational expression does not have a degree, so i have absolutely no idea what this means

OpenStudy (epicteatime):

Exactly! Thats why i was like, with

OpenStudy (epicteatime):

@satellite73 i need to ask a simple question, totally irrelevant from this topic

OpenStudy (epicteatime):

\[x ^{2}+x ^{2} is 2x ^{2} \right?\]

OpenStudy (epicteatime):

\[x ^{2} + x ^{2} is 2x ^{2}?\]

OpenStudy (epicteatime):

nvm i got it